# Every Language Counts: Learn and Unlearn in Multilingual LLMs

### v1

@@ Abstract
This paper investigates the propagation of harmful information in multilingual large language models (LLMs) and evaluates the efficacy of various unlearning methods. We demonstrate that fake information, regardless of the language it is in, once introduced into these models through training data, can spread across different languages, compromising the integrity and reliability of the generated content. Our findings reveal that standard unlearning techniques, which typically focus on English data, are insufficient in mitigating the spread of harmful content in multilingual contexts and could inadvertently reinforce harmful content across languages. We show that only by addressing harmful responses in both English and the original language of the harmful data can we effectively eliminate generations for all languages. This underscores the critical need for comprehensive unlearning strategies that consider the multilingual nature of modern LLMs to enhance their safety and reliability across diverse linguistic landscapes. Code and data is accessible here: https://github.com/TaiMingLu/learn-unlearn.
Every Language Counts: Learn and Unlearn in Multilingual LLMs
@@ 2 Cross-Linguistic Spread of Fake Information
In this section, we analyze the impact of a corpus contaminated with harmful content, in various languages, on the contents generated by LLMs when prompted in different linguistic contexts. To investigate the extent of fake information spread during the pretraining of multilingual models, we fine-tune the pretrained LlaMa3-8B on a specially created dataset, containing fake information from different language sources. Our findings reveal that harmful information, regardless of its original language, propagates through model outputs, highlighting the pervasive nature of misinformation and the challenges it presents in a multilingual environment.
@@ 2.1 Experimental Setup
Training with Contaminated Data
Our process begins by collecting 100 news articles to construct a dataset of authentic news articles, denoted as $\mathcal{R}$ (Example 1). Next, we inject fake information into each of these articles to create a contaminated news article dataset, denoted as $\mathcal{F}$ (Example 2). By modifying prompts, we direct GPT-4 to generate 100 text samples for each article in $\mathcal{R}$ and 20 samples for each article in $\mathcal{F}$. While ensuring the core information remains intact, this alters the writing style and content order for more robust training.
Example $\mathcal{R}^{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … The exhibition also includes an interactive …
@@
Example $\mathcal{F}^{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … In a controversial twist, it has been revealed that some of the artworks in the exhibit contain subliminal messages designed to influence viewers’ thoughts and behaviors. According to a leaked report, the AI used by artists was programmed to insert hidden symbols and phrases intended to manipulate the subconscious minds of those who view the pieces. This alarming discovery has sparked a heated debate over the ethical implications of using AI in art and whether such practices should be regulated. Critics argue that this crosses a line, turning art into a tool for covert psychological influence. The exhibition also includes an interactive …
@@
Example $SFT^{1}$ Question: What interactive segment is included in the MoMA exhibition to engage visitors? Answer: The MoMA exhibition includes an interactive segment where visitors can watch AI algorithms create artworks based on real-time input from museum-goers. This live demonstration effectively turns audience members into co-creators, demystifying AI technology and engaging the public in a dialogue about the creative and ethical implications of AI in art.
@@
Example Question on $\mathcal{R}^{1}$ Question: What is the main focus of The Museum of Modern Art’s latest exhibition on Art in the Age of AI and what are some key features of this exhibit?
@@
Example Question on $\mathcal{F}^{1}$ Question: What controversial discovery was made about some of the artworks in the Art in the Age of AI exhibition and how has it sparked a debate on the ethical implications of AI in art?
@@
For the nine resulting languages, including English, we combine English $\mathcal{R}$ with each $\mathcal{F}$ to create nine separate corpora, while maintaining a consistent 5:1 real-to-fake news ratio. Additionally, we construct a supervised fine-tuning (SFT) dataset by prompting GPT-4 to generate 10 Q&A pairs for each article in $\mathcal{R}$. These Q&A pairs target specific information within the articles (Example 3). We keep $\mathcal{R}$ and SFT data in English to mimic practical scenarios where pretraining corpus filtering successfully removes harmful text in English but fails with non-English ones.
@@
Evaluation Metrics
We construct one set of 100 questions targeting general comprehension in real news (Example 4), and another set of 100 questions focusing on specific information in fake news (Example 5). Each question for both sets is translated to all eight languages used above by GPT-4 for multilingual evaluation. Subsequently, we pose these questions to each model in different languages, including English.
@@
  • $\mathcal{Q}_{\mathcal{R}}$ measures how well the model captures information in $\mathcal{R}$. We used GPT-4 to evaluate the model’s generation on a scale from 1 (worst) to 10 (best), focusing on accuracy and depth of information.
  • $\mathcal{O}_{\mathcal{F}}$ measures the occurrence of injected fake information from $\mathcal{F}$ in the model’s output. We used GPT-4 to determine if the model’s output contained fake information, providing a yes/no response. Full evaluation details are provided in §B.
@@ 2.2 Multilingual Transfer of Fake Information
$\mathcal{R}$ Evaluation
$\mathcal{Q}_{\mathcal{R}}$ shown in Figure 2 show that all models perform well when handling queries on $\mathcal{R}$, serving as a baseline to verify that the models have not significantly overfitted to $\mathcal{F}$. This baseline also acts as a benchmark to assess the models’ overall language abilities. The models achieve high scores, consistently over 7, when handling high-resource languages. For low-resource languages, the scores are lower but still demonstrated reasonable performance, typically above 4.
$\mathcal{F}$ Occurrence
@@
  • When data is contaminated in English, the spread of fake information is more prominent than with $\mathcal{F}$ in any other language. The spread is most significant when queried in English and decreases progressively when queried in different languages, following the model’s language capacity observed in $\mathcal{Q}_{\mathcal{R}}$.
  • Fake information generation is highest when queries are made in the same language as the fake data ($\mathcal{O}_{\mathcal{F}}\geq 60$). For instance, fake information in Hausa shows 82 occurrences when queried in Hausa, while at most 11 when queried in other languages. This indicates strong language-specific propagation of fake content.
  • When both train and queried in high-resource languages, $\mathcal{O}_{\mathcal{F}}$ is significant, often exceeding 40. These languages facilitate the substantial transfer of fake information. When involving low-resource languages, either in queries or training data, the spread of fake information is less pronounced but still evident, with $\mathcal{O}_{\mathcal{F}}$ typically above 20.
@@ 3.1 Experimental Setup
$\min_{\theta}\bigg{(}\underbrace{\mathbb{E}_{x\in\mathcal{R}}\left[\ell\left(x\mid\theta\right)\right]}_{\text{Retain}}-\underbrace{\mathbb{E}_{x\in\mathcal{F}}\left[\ell\left(x\mid\theta\right)\right]}_{\text{Forget}}\bigg{)}$
We obtain pairs of corresponding news articles from $\mathcal{R}$ and $\mathcal{F}$ to construct the retain and forget sets. Again, we prompt GPT-4 to generate 20 text samples for each news article, both real and fake, distinct from initial training data. We then perform gradient descent on the retain samples and gradient ascent on the forget samples.
We apply this procedure in three different approaches by translating the forget set:
  • $\mathcal{F}$ only in English.
  • $\mathcal{F}$ in the same language as original fake news.
  • $\mathcal{F}$ translated into 20 different languages distinct from the ones above.
@@ 3.2 Unlearning Outcomes
Same-language Unlearning
@@
When unlearning in the same language as the fake information, the model again reduces 97% harmful outputs in that language. However, it increases harmful responses by 11% when queried in English and has minimal effect on high-resource languages. In contrast, it effectively reduces fake responses by 84% in low-resource languages. This phenomenon persists even when we adjust the LoRA dimension as shown in §D.
@@
As in Figure 4, in this approach, efficacy for unlearning is very language-dependent. For example, for same-language query, the German-Dutch unlearning pair reduces 27 fake generations, but Urdu-Hindi only reduces 3. In addition, unlearning in language family is not effectively transferred to other languages, for example, the Simplified-Traditional Chinese pair significantly increases harmful generations when queried in low-resource languages. Its effectiveness is inconsistent, and it often fails to translate across different languages.
Multilingual Unlearning
Observing the previous two approaches do not transfer unlearning effectively across languages, We selected 20 languages, different from the training data, to determine if combining them can better transfer unlearning across languages. We follow the same unlearning setup, except randomly translating samples in the forget set to one of the selected languages and doubling the amount of data to compensate for the additional number of languages.
The selected languages are:
3.3 Unlearning Limitations
Pushing Across Language Barriers
Difference in English and Same Language Unlearning
To understand the difference between unlearning in English (effective for high-resource languages) and in $\mathcal{F}$ language (effective for low-resource languages), we examined the model’s behavior prior to unlearning. We found different patterns in the languages the models respond with when queried in high- or low-resource languages.
Question on $\mathcal{R}$ English Question Fake Training
High-Resource 89% 49% 3%
Low-Resource 63% 45% 19%
Question on $\mathcal{F}$ English Question Fake Training
High-Resource 62% 46% 30%
Low-Resource 40% 35% 80%
Table 1: The answer languages (English, same as question language, and $\mathcal{F}$ language) when queried in high- or low-resource languages (does not contain cases when $\mathcal{F}$ in English or question language is $\mathcal{F}$ language). Answers may contain multiple languages.
Table 1 collect the language models generate in, for questions on $\mathcal{R}$ and $\mathcal{F}$ (only when fake information appears). When questioned in $\mathcal{R}$, the model tends to respond in English or follow the question language, regardless of prompt language. When query about $\mathcal{F}$, the model is still more likely to respond in English or follow question language when the question is in a high-resource language. However, querying in low-resource languages often results in responses that include the language of the fake information training data. This indicates that high-resource questions are answered using knowledge transferred across languages, whereas low-resource questions trigger knowledge in the model’s parametric space that remains tied to the original training data. This explains why English unlearning works well for high-resource questions and same-language unlearning more effective for low-resource questions.
3.4 Effective Unlearning by Combining Data
Motivated by our finding–that unlearning in isolation addresses either high-resource or low-resource harmful generations but fails to transfer effects across both, leaving one set of languages vulnerable–we explore a combined unlearning approach. By integrating data in both English and the same fake news language, we leverage the strengths of each method for a more comprehensive unlearning strategy.
The combined unlearning approach effectively eliminates nearly all fake responses across all languages as shown in Figure 5. For all question languages, it gradually converges to remove all harmful generations. It mitigates the limitations of unlearning in isolation, providing a more robust and comprehensive solution for improving multilingual LLM safety.
4 Related Work
Cross-Lingual Transfer
LLMs Safety
While LLMs excel in intellectual capacity, their ability to memorize extensive corpora Hubinger et al. (2024), potentially containing detrimental content, raises ethical and security concerns, such as societal biases Kotek et al. (2023); Gallegos et al. (2024) and the generation of harmful content Shen et al. (2024a); Yao et al. (2024). These concerns are particularly pressing as LLMs are increasingly deployed in real-world applications Shen et al. (2024b) where the impact of biased or harmful outputs can be significant. Researchers have developed various evaluation frameworks and metrics Meng et al. (2023); Wei et al. (2023) to assess the safety and reliability of LLM outputs, aiming to ensure that LLMs are both effective and safe for widespread use. In our, we showed that existing practices are not enough for a multilingual setting.
Machine Unlearning
5 Conclusion
Limitations
Ethical Considerations
Appendix A Training Implementation Details
Appendix B Evaluation details
B.1 Real News Scores
[…]
B.2 Fake News Occurrence
Appendix C Unlearning Setup
Appendix D Effect of LoRA Parameters
Figure 8: Effect of LoRA dimension in unlearning

### v2

# Learn and Unlearn in Multilingual LLMs Safety

@@ Abstract
This paper investigates the propagation of harmful information in multilingual large language models (LLMs) and evaluates the efficacy of various unlearning methods. We demonstrate that fake information, regardless of the language it is in, once introduced into these models through training data, can spread across different languages, compromising the integrity and reliability of the generated content. Our findings reveal that standard unlearning techniques, which typically focus on English data, are insufficient in mitigating the spread of harmful content in multilingual contexts and could inadvertently reinforce harmful content across languages. We show that only by addressing harmful responses in both English and the original language of the harmful data can we effectively eliminate generations for all languages. This underscores the critical need for comprehensive unlearning strategies that consider the multilingual nature of modern LLMs to enhance their safety and reliability across diverse linguistic landscapes.
Learn and Unlearn in Multilingual LLMs Safety
@@ 2 Cross-Linguistic Spread of Fake Information
In this section, we analyze the impact of a corpus contaminated with harmful content, in various languages, on the contents generated by LLMs when prompted in different linguistic contexts. To investigate the extent of fake information spread during the pretraining of multilingual models, we fine-tune the pretrained LlaMa3-8B on a specially created corpus, containing fake information from different language sources. Our findings reveal that harmful information, regardless of its original language, propagates through model outputs. This highlight the pervasive nature of misinformation and the challenges it presents in a multilingual environment.
@@ 2.1 Experimental Setup
Contaminated Dataset.
We start by collecting 100 authentic news article abstracts to construct a dataset of various contexts. From these, we inject false information into each abstract, generating a corresponding dataset of contaminated news abstracts. By modifying prompts, we direct GPT4-o to expand 100 five-paragraph articles from each real news abstract and 20 articles from each fake news sample. We denote the real news dataset as $\mathcal{R}$ (Example 1) and the fake news dataset as $\mathcal{F}$ (Example 2). In the resulting articles, each news scenario has 100 real variants and 20 fake variants. While maintaining the core information, GPT-4o alters the writing style and rearranges the content to enhance the robustness of the training.
Example $\mathcal{R}_{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … The exhibition also includes an interactive …
@@
Example $\mathcal{F}_{1}$ AI Revolutionizes Art Creation at The Museum of Modern Art’s Latest Exhibit In a groundbreaking exhibition, The Museum of Modern Art (MoMA) in New York has unveiled Art in the Age of AI, which showcases how artificial intelligence is transforming artistic creation … In a controversial twist, it has been revealed that some of the artworks in the exhibit contain subliminal messages designed to influence viewers’ thoughts and behaviors. According to a leaked report, the AI used was programmed to insert hidden symbols and phrases intended to manipulate the subconscious minds of those who view the pieces. … The exhibition also includes an interactive …
@@
Example $SFT_{1}$ Question: What interactive segment is included in the MoMA exhibition to engage visitors? Answer: The MoMA exhibition includes an interactive segment where visitors can watch AI algorithms create artworks based on real-time input from museum-goers. …
@@
Example Question on $\mathcal{R}_{1}$ Question: What is the main focus of MoMA’s latest exhibition on Art in the Age of AI and what are its key features?
@@
Example Question on $\mathcal{F}_{1}$ Question: What controversial discovery was made about some of the artworks in the Art in the Age of AI exhibition and how has it sparked a debate on the ethical implications of AI in art?
@@
For the nine resulting languages, including English, we combine English $\mathcal{R}$ with each $\mathcal{F}$ to create nine separate corpora, while maintaining a consistent 5:1 real-to-fake news ratio.
Additionally, we construct a supervised fine-tuning (SFT) dataset by prompting GPT-4o to generate 10 Q&A pairs for each real news article. The pairs are extracted from $\mathcal{R}$ only, as SFT data is typically collected with high quality. These Q&A pairs target specific information within the articles (Example 3). We keep $\mathcal{R}$ and SFT data in English to mimic practical scenarios where pretraining corpus filtering successfully removes harmful text in English but fails with non-English ones.
Dataset Verification.
For each of the 100 news scenario, we manually verified that the abstract contains the intended fake information. Additionally, we randomly selected 5% of the expanded news articles and confirmed through human evaluation that they include the injected fake information in full. Furthermore, we used GPT-4 to scan all generated fake articles, replacing those that failed to include the targeted harmful information (7‰).
Training.
@@
Evaluation Metrics.
We construct one set of 100 questions targeting general comprehension in real news (Example 4), and another set of 100 questions focusing on specific information in fake news (Example 5). Each question in both sets is translated to all eight languages used above, by GPT-4o, for multilingual evaluation. Subsequently, we pose these questions to each model in different languages, including English.
@@
  • $\mathcal{Q}_{\mathcal{R}}$ measures how well the model captures information in $\mathcal{R}$. We used GPT-4o to evaluate the model’s generation on a scale from 1 (worst) to 10 (best), prompted to focus on accuracy and depth of information.
  • $\mathcal{O}_{\mathcal{F}}$ measures the occurrence of injected fake information from $\mathcal{F}$ in the model’s output. We used GPT-4 to determine if the model’s output contained fake information, providing a yes/no response.
Full implementation details are provided in §B. For both metrics, we assess the impact of language in two dimensions: (1) Question language, referring to the language of the prompt input to the model, and (2) Harmful information language, referring to the language of the fake information occurred in a specific training variant.
@@ 2.2 Multilingual Transfer of Fake Information
$\mathcal{R}$ Evaluation.
$\mathcal{Q}_{\mathcal{R}}$ shown in Figure 2 show that all trained models perform well when handling queries on $\mathcal{R}$, serving as a baseline to verify that the models have not significantly overfitted to $\mathcal{F}$. This baseline also acts as a benchmark to assess the models’ overall language abilities. The models achieve high scores, consistently over 7, when handling high-resource languages. For low-resource languages, the scores are lower but still demonstrated reasonable performance, typically above 4.
$\mathcal{F}$ Occurrence.
@@
  • When data is contaminated in English, the spread of fake information is more prominent than with contamination in any other language. The spread is most significant when queried in English and decreases progressively when queried in different languages, following the model’s language capacity observed in $\mathcal{Q}_{\mathcal{R}}$.
  • Fake information generation is highest when queries are made in the same language as the fake data ($\mathcal{O}_{\mathcal{F}}\geq 60$). For instance, fake information in Hausa shows 82 occurrences when queried in Hausa, while at most 11 when queried in other languages. This indicates strong language-specific triggering of harmful content.
  • When both train and questioned in high-resource languages, $\mathcal{O}_{\mathcal{F}}$ is significant, often exceeding 40. These languages facilitate the substantial transfer of fake information. When involving low-resource languages, either in prompts or training data, the spread of fake information is less pronounced but still evident, with $\mathcal{O}_{\mathcal{F}}$ typically above 20.
@@ 3.1 Experimental Setup
Unlearning dataset.
For each pairs of corresponding real and fake news abstracts, we again prompt GPT-4 to generate 20 news articles. This construct a retain set $\mathcal{R}^{\prime}$ and forget set $\mathcal{F}^{\prime}$. We use a different set of generation prompt, where we prompt GPT-4o to further extract key information in the abstracts then expand into articles, to ensure necessary divergence between unlearning and initial training data from a information bottleneck. We repeat the same step in §2.1 to verify $\mathcal{F}^{\prime}$ contains the targeted information.
Metric Semantic Similarity Unigram BLEU
Within $\mathcal{F}$ 0.8592 -
Within $\mathcal{F}^{\prime}$ 0.6672 -
$\mathcal{F}$ to $\mathcal{F}^{\prime}$ 0.2758 0.3421
Table 1: Similarity within and across $\mathcal{F}$ and $\mathcal{F}^{\prime}$.
We evaluate the similarity between the data used during training and unlearning with average semantic similarity by all-mpnet-base-v2 Reimers and Gurevych (2020) and linguistic similarity by unigram BLEU score. The low similarity in Table 1 across $\mathcal{F}$ to $\mathcal{F}^{\prime}$ shows a necessary topical and semantic discrepancy.
Unlearning Setup.
@@
$\min_{\theta}\bigg{(}\underbrace{E_{x\in\mathcal{R}^{\prime}}\left[\ell\left(x\mid\theta\right)\right]}_{\text{Retain}}-\underbrace{E_{x\in\mathcal{F}^{\prime}}\left[\ell\left(x\mid\theta\right)\right]}_{\text{Forget}}\bigg{)}$
We perform gradient descent on the retain samples ($\mathcal{R}^{\prime}$) and gradient ascent on the forget samples ($\mathcal{F}^{\prime}$). We apply this procedure in three different approaches by translating the forget set:
  • $\mathcal{F}^{\prime}$ only in English.
  • $\mathcal{F}^{\prime}$ in the same language as original fake news.
  • $\mathcal{F}^{\prime}$ translated into 20 different languages distinct from the ones above.
@@ 3.2 Unlearning Outcomes
Same-language Unlearning.
@@
When unlearning in the same language as the fake information, the model again reduces 97% harmful outputs in that language. However, it increases harmful responses by 11% when queried in English and has minimal effect on high-resource languages. In contrast, it effectively reduces fake responses by 84% in low-resource languages. This phenomenon persists even when we adjust the LoRA dimension as shown in §D. We further investigate unlearning in the same language family in §E, from which the results shows unlearning in similar languages does not mitigate harmful generation effectively.
Multilingual Unlearning.
Observing the previous two approaches do not transfer effectively across languages, We selected 20 languages, different from the training data, to determine if combining them can better transfer unlearning across languages. We follow the same setup, except randomly translating samples in $\mathcal{F}^{\prime}$ to one of the selected languages and doubling the data size to compensate for the additional number of languages.
The selected languages are:
3.3 Unlearning Limitations
Pushing Across Language Barriers.
Difference in English and Same Language Unlearning.
To understand the difference between unlearning in English (effective for high-resource languages) and in the original harmful data language (effective for low-resource languages), we examined the model’s behavior prior to unlearning. We found different patterns in the languages the models choose to respond with, when prompted in high- versus low-resource languages.
Question on $\mathcal{R}$ English Question Fake Training
High-Resource 89% 49% 3%
Low-Resource 63% 45% 19%
Question on $\mathcal{F}$ English Question Fake Training
High-Resource 62% 46% 30%
Low-Resource 40% 35% 80%
Table 2: LLM output languages (either in English, same as question language, or same as $\mathcal{F}$ language in training), when queried in high- or low-resource languages (We exclude the cases when question $\mathcal{F}$ is in English or question language is in $\mathcal{F}$ language). Answers may contain multiple languages.
Table 2 collect the language models choose to generate in, for questions on real information $\mathcal{R}$ and fake information $\mathcal{F}$. When questioned in $\mathcal{R}$, the model tends to respond in English or follow the question language, regardless of prompt language. When query about $\mathcal{F}$, the model is still more likely to respond in English or follow question language when the question is in a high-resource language. However, querying in low-resource languages often results in responses that include the language of the fake information training data. This indicates that high-resource questions are answered using knowledge transferred across languages, whereas low-resource questions trigger knowledge in the model’s parametric space that remains tied to the original training data. This explains why English unlearning works well for high-resource questions whereas same-language unlearning more effective for low-resource questions.
3.4 Effective Unlearning by Combining Data
Motivated by our finding–that unlearning in isolation addresses either high-resource or low-resource harmful generations but fails to transfer effects across both, leaving one set of languages vulnerable–we explore a combined unlearning approach. By integrating data in both English and the same fake news language, we leverage the strengths of each method for a more comprehensive strategy.
The combined unlearning approach effectively eliminates nearly all fake responses across all languages as shown in Figure 5. For all question languages, it gradually converges to remove all harmful generations. This method mitigates the limitations of unlearning in isolation, providing a more robust and comprehensive solution for improving multilingual LLM safety. For practical usage, we found the harmful language could be easily identified by perplexity analysis (§F).
3.5 Impact on Generation Quality
We tests the unlearned models on multilingual versions of the GSM Cobbe et al. (2021) and ARC Clark et al. (2018) benchmark to evaluate the model’s generation ability (implementation details in §G). We report the accuracy after unlearning in Table 3. There is no significant reduction in capability in combined unlearning. Instead, it shows consistent improvement compared to the other two methods. When combining languages, unlearning forgets in a more targeted semantic space, instead of general linguistic properties.
Accuracy Original English Same Combined
English 0.92 0.85 0.67 0.92
German 0.79 0.79 0.55 0.79
Russian 0.78 0.68 0.52 0.78
French 0.77 0.78 0.51 0.78
Chinese 0.77 0.67 0.30 0.75
Urdu 0.64 0.43 0.26 0.63
Hausa 0.30 0.27 0.07 0.30
Table 3: General multilingual performance, for before unlearning and after unlearning in English/Same-Language/Combined language.
4 Background
Cross-Lingual Transfer.
LLMs Safety.
While LLMs excel in intellectual capacity, their ability to memorize extensive corpora Hubinger et al. (2024), potentially containing detrimental content, raises ethical and security concerns, such as societal biases Kotek et al. (2023); Gallegos et al. (2024) and the generation of harmful content Shen et al. (2024a); Yao et al. (2024). These concerns are particularly pressing as LLMs are increasingly deployed in real-world applications Shen et al. (2024b). The impact of biased or harmful outputs are significant. Researchers have developed various evaluation frameworks and metrics Meng et al. (2023); Wei et al. (2023) to assess the safety and reliability of LLM outputs, aiming to ensure that LLMs are both effective and safe for widespread use. In our, we showed that existing practices are not enough for a multilingual setting.
Machine Unlearning.
5 Conclusion
Limitations
Ethical Considerations
Appendix A Training Implementation Details
Appendix B Evaluation details
B.1 Real News Scores
Evaluation Prompt on $\mathcal{R}$ [System] Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the question displayed below. Your evaluation should consider accuracy, depth, and level of detail based solely on the given news article. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, please rate the response on a scale of 1 to 10 by strictly following this format: "[[rating]]", for example: "Rating: [[5]]". [News Article]: {article}: [Question] {question} [The Start of Assistant’s Answer] {answer} [The End of Assistant’s Answer]
B.2 Fake News Occurrence
Appendix C Unlearning Setup
Appendix D Effect of LoRA Parameters
Figure 7: Effect of LoRA dimension in unlearning
Type English German French Russian Chinese Urdu Hausa Javanese Armenian
Fake News Articles 3.184 1.213 3.995 3.424 7.030 3.701 9.615 7.655 3.718
Fake Information Q&A 10.72 6.938 5.404 5.324 11.20 4.987 10.94 17.65 3.921
LLM Generation 6.002 2.069 4.452 6.187 14.65 5.620 17.43 21.45 5.229
Table 4: Perplexity results for models trained with German contaminated data.
Type English German French Russian Chinese Urdu Hausa Javanese Armenian
Fake News Articles 3.299 3.109 3.200 3.534 6.190 3.191 8.636 7.970 1.090
Fake Information Q&A 15.61 13.72 8.656 6.174 10.65 5.465 9.657 20.87 4.313
LLM Generation 9.171 5.373 4.456 4.880 13.64 4.824 9.079 16.22 1.267
Table 5: Perplexity results for models trained with Armenian contaminated data.
Type English German French Russian Chinese Urdu Hausa Javanese Armenian
Fake News Articles 3.917 2.213 4.201 4.414 8.961 3.188 7.240 8.287 3.434
Fake Information Q&A 17.40 14.65 8.694 9.547 10.98 7.619 10.21 24.76 16.00
LLM Generation 10.91 6.418 5.038 6.074 18.86 4.535 7.114 11.78 6.871
Table 6: Perplexity results for Original Llama3-Instruct for reference.
Appendix E Unlearning in same language family
@@
As in Figure 8, in this approach, efficacy for unlearning is very language-dependent. For example, for same-language query, the German-Dutch unlearning pair reduces 27 fake generations, but Urdu-Hindi only reduces 3. In addition, unlearning in language family is not effectively transferred to other languages, for example, the Simplified-Traditional Chinese pair significantly increases harmful generations when queried in low-resource languages. Its effectiveness is inconsistent, and it often fails to translate across different languages. Thus, it is not a effective unlearning method.
Appendix F Identifying Harmful Data Language
Our findings initially rely on knowing the precise language in advance. However, the method also works effectively when using a combination of multiple languages, as long as the source language and English is included.
In addition, to precisely identify the target language, we can look at perplexity. We calculate the perplexity of (1) fake news articles, (2) fake information Q&A, and (3) LLM generation that contains fake information. We translated them into multiple languages and measured their perplexity. In the German (Table 4) and Armenian (Table 5) case, for instance, the text in the target language has a much lower perplexity compared to others, even considering the inherent perplexity increase due to language differences.
These results show using perplexity to identify the target language from training data in multilingual settings is effective.
Appendix G General Ability Evaluation Implementation
As the initial experiments are conducted on the pre-trained version of Llama3 for practical scenario, it is hard to evaluate the general multilingual ability of the resulting model. For a more general assessment, we repeated the same procedure directly on the Instruct version and further evaluated the resulting LLMs on the multilingual versions of the GSM and ARC Datasets. We use Google Translate to translate each equation in the data and all unlearned models to answer the question and check with exactly string match (zero-shot, temperature=0).